The displacement ''x'' and time of travel ''t'' for a particle moving on a straight line are related as t=sqare root (x+1)(x-1). Its acceleration at time ''t'' is?Please show the solution.
a)1/x-1/xsquare
b)1/xcube
c)-tsquare/xcube
d)-t/xsquare

Grade:11

2 Answers

Aman Bansal
11 years ago

Dear Nimisha,

A 4 kg ball on a string is rotated about a circle of radius 10 m. The maximum tension allowed in the string is 52 N. What is the maximum speed of the ball?
Solution:

 
We have F = ma

where F = 52N, and m= 4 kg

a = [Math Processing Error] = [Math Processing Error] = 13m/s2
    
      And a =[Math Processing Error]

V = [Math Processing Error]
    =[Math Processing Error] 
    = 11.4 m/s.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

vihang agarwal
11 years ago

=>  t^2=x^2-1

=> x=sqr rt t^2+1

=> differentiate it twice with respect to time to get the accelaration

=> the ans will be  [(t^2+1) - t^2] / [(t^2+1)(sqr rt t^2+1)]

=> now substitute t^2 = x^2-1 to get acc = 1/x^3